OpenStudy (anonymous):

y=2x+7 Is the equation of a line. Explain how you can find two sets of (x, y) points that are on the line.

OpenStudy (texaschic101):

well....just sub in numbers. y = 2x + 7 ...pick a number for y...lets pick 3. Now sub in 3 for y. 3 = 2x + 7 -- subtract 7 from both sides 3 - 7 = 2x -4 = 2x -- divide by 2 -2 = x so one point on the line is (-2,3) understand how I did that ?

OpenStudy (texaschic101):

so basically, just pick a number for one of the variables, sub that into the equation, and then find the other variable.

OpenStudy (anonymous):

ohhh okayy i understand

OpenStudy (texaschic101):

cool...thats always good to hear :)

OpenStudy (anonymous):

thanks your the best

OpenStudy (texaschic101):

I don't know about that, but I am happy to help :)